OpenStudy (anonymous):

Akeem invests $25,000 in an account that pays 4.75% annual interest compounded continuously. Using the formula A=Pe^rt, where A=the amount in the account after t years, P=principal invested, and r=the annual interest rate, how many years, to the nearest tenth will it take Akeem’s investment to triple? SHOW WORK! a) 10.0 b) 14.6 c) 23.1 d) 24.0

OpenStudy (anonymous):

the normal equation is A=Pe^rt but now you want to find how long it'll take for your initial investment (P) to triple. so in this case the new tweaked equation would be 3P=Pe^rt from here all you do is just try to isolate the t Divide both sides by P and you get 3=e^rt If I'm not mistaken getting the "ln"of both sides would would reduce the e on the right side to 1 since ln(e)=1 leaving you with ln3=rt and then all you do is divide both sides by r and get ln(3)/ r= t which remember to have r in decimal format.

OpenStudy (anonymous):

if you have questions about the answer or something, let me know